# ResBit: Residual Bit Vectors for Categorical Values

### v3

@@ Abstract
One-hot vectors, a method for representing discrete/categorical data, are commonly used in machine learning due to their simplicity and intuitiveness. However, the one-hot vectors suffer from a linear increase in dimensionality, posing computational and memory challenges, especially when dealing with datasets containing numerous categories. To address this issue, we propose Residual Bit Vectors (ResBit), a technique for densely representing categorical data. While Analog Bits (Chen et al., 2023) presents a similar approach, it faces challenges in categorical data generation tasks. ResBit overcomes these limitations, offering a more versatile solution. In our experiments, we focus on tabular data generation, examining the performance across scenarios with varying amounts of categorical data. We verify the acceleration and ensure the maintenance or improvement of performance.
Machine Learning, ICML, tabular data generation, curse of dimensionality
@@ 1 Introduction
To address the dimensionality increase of one-hot vectors accompanying the growth in cardinality, we propose Residual Bit Vectors (ResBit) by integrating the concepts of Analog Bits (Chen et al., 2023) and Residual Vector Quantization (RVQ) (Juang & Gray, 1982) (Section 3.4). ResBit represents categorical data by hierarchically acquiring bit representations. Our research is motivated by the phenomenon of generation collapse observed in TabDDPM (Kotelnikov et al., 2023), one of the tabular data generative models. This phenomenon, akin to mode collapse, involves the generation of only specific values. This problem has also been encountered in other studies using TabDDPM as a baseline, as noted in the work of Zhang et al. (2023). While maintaining the same hyperparameter settings during the training of TabDDPM, we confirmed that the phenomenon of generation collapse is dependent on the cardinality of categorical data. This dependence is observed in conjunction with the dimensionality of the input to the multinomial diffusion models constituting TabDDPM, where the input is provided in the form of one-hot vectors. Based on this observation, we concluded that the dimensionality of one-hot vectors is a contributing factor to generation collapse. While conventional bit representations suffice for representing categorical data, they introduce specific problem in generation tasks (Section 3.3). ResBit addresses this problem while mitigating the increase in dimensionality even with growing cardinality (Section 3.5).
@@ 3.2 Generation Collapse in TabDDPM
For a theoretical explanation of this phenomenon, please refer to Appendix C.
@@ 3.4 Residual Bit Vectors
$M-1=\sum_{k=0}^{\infty}a_{k}(2^{k}-1)\quad\mathrm{s.t.}\min\sum_{k=0}^{\infty}a_{k}k$ (2)
@@
$N=\sum_{k=1}^{\infty}c_{k}(2^{k}-1)\quad(c_{k}\in[0,a_{k}],c_{k}\in\mathbb{Z})$ (3)
@@ 4.1 Experimental Setups
We use 10 datasets to conduct a comprehensive evaluation of the existing methods and verify the effectiveness of our proposed method. The basic overview of the dataset is presented in Table 2. For additional details such as the number of training data, please refer to Table 9 in Appendix A.4. ResBit is a method representing categorical data. Therefore, experiments are conducted on datasets that include both categorical and numerical data. To validate the effectiveness of our proposed method, we prepare datasets with both low and high cardinalities. Low cardinality datasets primarily involve datasets used in TabDDPM (Kotelnikov et al., 2023). There are two main purposes for evaluating on datasets of these two different scales. One purpose is to verify the performance of the proposed method on low cardinality datasets. The other is to assess the effectiveness of the proposed method and existing methods on high cardinality datasets. Most datasets are real-world datasets, except for the CC dataset, which is a synthetic dataset.
Table 2: Brief overview of the dataset used in our experiments.
@@ 5 Conclusion and Future Work
In this paper, we proposed ResBit as a method to address the “out-of-index” problem in generative tasks. ResBit acquires hierarchical bit representations, overcoming the “curse of dimensionality” faced by one-hot vectors, especially in the case of high cardinality categorical features. In our experiments, when applied to TabDDPM, which motivated our research, ResBit achieved a speedup in training and generation while maintaining accuracy. Additionally, it enabled the generation of diverse categorical values. Through comprehensive experiments, we highlighted shortcomings in existing tabular data generation models. Despite achieving high-speed sampling, training times were notably long, and these models did not demonstrate effective results in configurations not explored in previous researches.
In the future, further validation of ResBit is desired. For instance, although we confirmed its effectiveness in image classification and class-conditional image generation tasks in Appendix E, the experiments were not exhaustive, and there are other potential applications, such as image segmentation or algorithms utilizing one-hot vectors. Furthermore, as existing methods for tabular data generation all convert categorical data into one-hot vectors, constructing models with fewer parameters becomes possible by applying ResBit. While we demonstrated its application to TabDDPM in this paper, its effectiveness is likely to extend to other models as well.
6 Impact Statement
Our research holds the potential for significant dimensionality reduction compared to traditional one-hot vectors by utilizing dense vectors, thereby enabling possibilities such as parameter reduction in deep learning and efficient operations.
@@ Appendix B Details of the cardinalities survey
Appendix C Mutlinomial Diffusion Analysis from the Fomulation
We consider mode collapse from the perspective of the multinomial diffusion equation, the forward process of multinomial diffusion is formulated using a categorical distribution as follows:
$\lim_{K\to\infty}\boldsymbol{\pi}=[\alpha_{t}\boldsymbol{x}_{t}]\odot[\overline{\alpha}_{t-1}\boldsymbol{x}_{0}]=(\alpha_{t}x_{t,i}\cdot\overline{\alpha}_{t-1}x_{0,i})_{0\leq i\leq K}=\alpha_{t}x_{t,k}\cdot\overline{\alpha}_{t-1}$ (9)
Therefore,
Then,
@@ E.2 Class Conditional Image Generation
We conduct experiments using Conditional Generative Adversarial Networks (CGAN) (Mirza & Osindero, 2014) as a method to specify classes. In other words, we concatenate a noise vector and a conditional vector representing the class and input them to the Generator. We use the CIFAR-10 (Krizhevsky, 2009), MNIST (LeCun et al., 2010), and Food101 datasets. For the architectures, we utilize InfoGAN (Chen et al., 2016) for CIFAR-10 and MNIST, and SNGAN (Miyato et al., 2018) for Food101. We use Adam (Kingma & Ba, 2017) as the optimizer with hyperparameters $lr=0.0002,\beta_{1}=0.5,\beta_{2}=0.999$. Our primary objective is not to improve the quality of generated images, so we do not incorporate accuracy improvement techniques. However, in the case of SNGAN, we aim for training stability by updating the discriminator five times for every single update of the generator. We conduct training for 30 epochs on MNIST, 300 epochs on CIFAR-10, and updated the generator 15,000 times for Food101. Results are presented in Figures 6, 7, and 8. It is evident that conditioning works for all datasets. In MNIST, ResBit lacks diversity, while in CIFAR-10, visually, it can be observed that one-hot has more mode collapse.

### v4

@@ Abstract
One-hot vectors, a common method for representing discrete/categorical data, in machine learning are widely used because of their simplicity and intuitiveness. However, one-hot vectors suffer from a linear increase in dimensionality, posing computational and memory challenges, especially when dealing with datasets containing numerous categories. In this paper, we focus on tabular data generation, and reveal the multinomial diffusion faces the mode collapse phenomenon when the cardinality is high. Moreover, due to the limitations of one-hot vectors, the training phase takes time longer in such a situation. To address these issues, we propose Residual Bit Vectors (ResBit), a technique for densely representing categorical data. ResBit is an extension of analog bits and overcomes limitations of analog bits when applied to tabular data generation. Our experiments demonstrate that ResBit not only accelerates training but also maintains performance when compared with the situations before applying ResBit. Furthermore, our results indicate that many existing methods struggle with high-cardinality data, underscoring the need for lower-dimensional representations, such as ResBit and latent vectors.
Machine learning, tabular data generation, curse of dimensionality
@@ 1 Introduction
To address the dimensionality increase of one-hot vectors accompanying the growth in cardinality, we propose Residual Bit Vectors (ResBit) by integrating the concepts of Analog Bits (Chen et al., 2023) and Residual Vector Quantization (RVQ) (Juang & Gray, 1982) (Section 3.4). ResBit represents categorical data by hierarchically acquiring bit representations. Our research is motivated by the phenomenon of generation collapse observed in TabDDPM (Kotelnikov et al., 2023), one of the tabular data generative models. This phenomenon, akin to mode collapse, involves the generation of only specific values. This problem has also been encountered in other studies using TabDDPM as a baseline, as noted in the works of Zhang et al. (2023) and Shi et al. (2024). While maintaining the same hyperparameter settings during the training of TabDDPM, we confirmed that the phenomenon of generation collapse is dependent on the cardinality of categorical data. This dependence is observed in conjunction with the dimensionality of the input to the multinomial diffusion models constituting TabDDPM, where the input is provided in the form of one-hot vectors. Based on this observation, we concluded that the dimensionality of one-hot vectors is a contributing factor to generation collapse. While conventional bit representations suffice for representing categorical data, they introduce specific problem in generation tasks (Section 3.3). ResBit addresses this problem while mitigating the increase in dimensionality even with growing cardinality (Section 3.5).
@@ 3.2 Generation Collapse in TabDDPM
For more details, we discuss it using simple synthetic data in Appendix C.
In terms of the formulation of multinomial diffusion, the forward process of multinomial diffusion is formulated using a categorical distribution as follows:
$\lim_{K\to\infty}\boldsymbol{\pi}$ $=[\alpha_{t}\boldsymbol{x}_{t}]\odot[\overline{\alpha}_{t-1}\boldsymbol{x}_{0}]$
$=(\alpha_{t}x_{t,i}\cdot\overline{\alpha}_{t-1}x_{0,i})_{0\leq i\leq K}=\alpha_{t}x_{t,k}\cdot\overline{\alpha}_{t-1}$ (5)
Therefore,
Then,
@@ 3.4 Residual Bit Vectors
$\min\sum_{k=0}^{\infty}a_{k}k\quad\mathrm{s.t.}\quad M-1=\sum_{k=0}^{\infty}a_{k}(2^{k}-1)$ (9)
@@
$N=\sum_{k=1}^{\infty}c_{k}\quad(c_{k}\in[0,a_{k}(2^{k}-1)],c_{k}\in\mathbb{Z})$ (10)
@@ 4.1 Experimental Setups
We use 10 datasets to conduct a comprehensive evaluation of the existing methods and verify the effectiveness of our proposed method. The basic overview of the dataset is presented in Table 2. For additional details such as the number of training data, please refer to Table 9 in Appendix A.4. ResBit is a method representing categorical data. Therefore, experiments are conducted on datasets that include both categorical and numerical data. To validate the effectiveness of our proposed method, we prepare datasets with both low and high cardinalities. Low cardinality datasets primarily involve datasets used in TabDDPM (Kotelnikov et al., 2023). There are three main purposes for evaluating on datasets of these two different scales.
  1. 1.
    Comparing the performance of the existing methods as a comprehensive evaluation
  2. 2.
    Verifying the performance of our proposed method on low cardinality datasets (maintaining the performance of TabDDPM)
  3. 3.
    Assessing the effectiveness of our proposed method and confirming the performance of the existing methods on high cardinality datasets
Most datasets are real-world datasets, except for the CC dataset, which is a synthetic dataset.
Table 2: Brief overview of the datasets used in our experiments.
@@ 5 Conclusion and Future Work
In this paper, we focus on the mode collapse phenomenon in TabDDPM, investigate its cause, and propose ResBit, a hierarchical bit representation, as a solution to avoid it. ResBit is an extension of analog bits to address the “out-of-index” problem. This hierarchical bit representation theoretically avoids the “out-of-index” problem. By treating ResBit as numerical data when applying it to tabular data generative models, training time is shortened because of the reduction in the dimensionality of categorial data, the sampling quality is maintained, and diverse categorical values are generated. Throughout our comprehensive experiments, we also observe that a multitude of existing methods do not demonstrate the effective results in the settings not explored in previous research such as high cardinality dataset and the large-scale dataset.
Although we focus on cardinalities, we believe there are other important factors, which have not received sufficient attention in tabular data generation. In addition, since the results are dependent on the dataset, further validation with a wider range of data is also required.
@@ Appendix B Details of the cardinalities survey
Appendix C Generation Collapse in TabDDPM using Simple Synthetic Data
We investigate mode collapse phenomenon in TabDDPM using simple synthetic data. We make data by fixing the numerical and target variables and changing the cardinalities of the categorical variable. The process of the way to make data is shown below.
  1. 1.
    Set $N=10000$, the number of rows.
  2. 2.
    For a numerical column, we set it to normal distribution.
  3. 3.
    For a target column, we set it to normal distribution.
  4. 4.
    For a categorical column, we use long tail distribution and uniform distribution respectively.
The cardinalities we use are 50, 100, 2000, and 5000. Therefore, the number of data types is eight. We train TabDDPM using the data and generate 10000 samples. Figure 6 shows the comparison of the distributions between original and generated data each cardinality. We observed only specific categorical values have appeared as the cardinalities increased regardless of what the categorical distribution is.
Uniform distribution Long tail distribution
50 cardinalities
Refer to caption Refer to caption
100 cardinalities
Refer to caption Refer to caption
2000 cardinalities
Refer to caption Refer to caption
5000 cardinalities
Refer to caption Refer to caption
Figure 6: Visualizations of original and generated distributions.
Moreover, we observe the categorical distributions in case the cardinality is 2000. We show the log scale visualized distributions to recognize the shape of distributions in Figure 7. When the categorical distribution is uniform, the model is not learning the distribution at all. On the other hand, when the categorical distribution is long tail, the generated distribution is similar that of original. This is because even if the model generates only certain classes, the shape of the distribution will be like a long tail. Therefore, it is possible to roughly learn the distribution if only the distribution of the top few classes in order of frequency can be learned. This is suggested by the fact that, although the tail part is not generated, the distribution of the top classes appears similar in Figure 7. However, when the categorical distribution is uniform, it is necessary for all classes to be generated equally. We consider that it is difficult to learn even the rough shape of the distribution due to this limitation.
Uniform distribution Long tail distribution
Refer to caption Refer to caption
Figure 7: Visualization of original and generated categorical distributions in log scale when the number of cardinality is 2000.
@@ E.2 Class Conditional Image Generation
We conduct experiments using Conditional Generative Adversarial Networks (CGAN) (Mirza & Osindero, 2014) as a method to specify classes. In other words, we concatenate a noise vector and a conditional vector representing the class and input them to the Generator. We use the CIFAR-10 (Krizhevsky, 2009), MNIST (LeCun et al., 2010), and Food101 datasets. For the architectures, we utilize InfoGAN (Chen et al., 2016) for CIFAR-10 and MNIST, and SNGAN (Miyato et al., 2018) for Food101. We use Adam (Kingma & Ba, 2017) as the optimizer with hyperparameters $lr=0.0002,\beta_{1}=0.5,\beta_{2}=0.999$. Our primary objective is not to improve the quality of generated images, so we do not incorporate accuracy improvement techniques. However, in the case of SNGAN, we aim for training stability by updating the discriminator five times for every single update of the generator. We conduct training for 30 epochs on MNIST, 300 epochs on CIFAR-10, and updated the generator 15,000 times for Food101. Results are presented in Figures 8, 9, and 10. It is evident that conditioning works for all datasets. In MNIST, ResBit lacks diversity, while in CIFAR-10, visually, it can be observed that one-hot has more mode collapse.